Chapter 14 Algorithmic Question 59 Part A The decomposition of N205 in solution in reaction 2N205 (soln) → 4NO2 (son) + 02 (soln) The reaction is first order and has a rate carbonechloride proceeds via the reaction is initiated with 0.055 mol in a 1.00-L vessel, how many moles remain after 151 s? O 0.027 M 2.1 x 103 M 0.052 M O 13 M 0.057 M Submit Previous Answers Request Answer
Answer #1

K   = 2.303/t log[A0]/[A]

[A0]   = 0.055mole

[A] =

t    = 151s

K   = 4.82*10^-3 s^-1

K   = 2.303/t log[A0]/[A]

4.82*10^-3 = 2.303/151 log0.055/[A]

log0.055/[A]    = 4.82*10^-3*151/2.303

log0.055/[A]   = 0.316

    0.055/[A]    = 10^0.316

   0.055/[A]    = 2.07

       [A]          = 0.055/2.07 = 0.027M >>>answer